# CViT: Continuous Vision Transformer for Operator Learning

### v2

@@ Abstract
Operator learning, which aims to approximate maps between infinite-dimensional function spaces, is an important area in scientific machine learning with applications across various physical domains. Here we introduce the Continuous Vision Transformer (CViT), a novel neural operator architecture that leverages advances in computer vision to address challenges in learning complex physical systems. CViT combines a vision transformer encoder, a novel grid-based coordinate embedding, and a query-wise cross-attention mechanism to effectively capture multi-scale dependencies. This design allows for flexible output representations and consistent evaluation at arbitrary resolutions. We demonstrate CViT’s effectiveness across a diverse range of partial differential equation (PDE) systems, including fluid dynamics, climate modeling, and reaction-diffusion processes. Our comprehensive experiments show that CViT achieves state-of-the-art performance on multiple benchmarks, often surpassing larger foundation models, even without extensive pretraining and roll-out fine-tuning. Taken together, CViT exhibits robust handling of discontinuous solutions, multi-scale features, and intricate spatio-temporal dynamics. Our contributions can be viewed as a significant step towards adapting advanced computer vision architectures for building more flexible and accurate machine learning models in the physical sciences.
@@ 2 Background and Related Work
The emergence of neural operators has sparked significant advancements in scientific machine learning, particularly in the realm of solving partial differential equations (PDEs) and modeling complex physical systems. Building upon the foundational architectures mentioned in the introduction, recent developments have focused on enhancing the expressiveness and efficiency of these models. For instance, extensions of the Fourier Neural Operator (FNO) (Li et al., 2021) have explored factorized representations (Tran et al., 2021) to reduce computational complexity while maintaining performance. Similarly, wavelet-based approaches (Gupta et al., 2021; Tripura & Chakraborty, 2022) have been proposed to capture multi-scale features more effectively. Most existing neural operator architectures can be interpreted as conditioned neural fields (Xie et al., 2022) with specific choices of base fields and conditioning mechanisms, see Appendices B and C for a detailed discussion.
Concurrently, the success of vision transformers in computer vision tasks has inspired their adaptation to scientific machine learning problems. The self-attention mechanism central to transformers offers a powerful tool for capturing long-range dependencies in spatial and temporal data, a crucial aspect in many PDE-governed systems. However, adapting vision transformers to the continuous nature of physical problems poses unique challenges, particularly in handling irregular domains and ensuring output consistency across different resolutions.
@@
In the specific context of operator learning, several transformer-based approaches have emerged to address these challenges. OFormer (Li et al., 2022) introduced a novel way to embed continuous input functions and output queries into a transformer architecture. This approach demonstrated the potential of transformers in handling the inherent continuity of physical problems. GNOT (Hao et al., 2023) further extended this idea by incorporating graph structures, allowing for more flexible representation of input functions and improved handling of irregular domains. A significant advancement in this direction came with DPOT (Hao et al., 2024), which introduced a denoising pre-training strategy. This approach enables the model to learn from diverse PDE datasets, significantly enhancing its generalization capabilities. However, these methods often require extensive computational resources for training and may struggle with maintaining consistency when evaluating at arbitrary resolutions.
@@
Despite these advancements, several key challenges remain in the field of operator learning. These include efficiently handling high-dimensional data while capturing long-range dependencies, ensuring output consistency and accuracy across different resolutions, and balancing model expressiveness with computational efficiency. Addressing even a subset of these challenges requires novel architectural designs that can leverage the strengths of both neural operators and vision transformers while overcoming their respective limitations. This sets the stage for our proposed Continuous Vision Transformer (CViT), which aims to tackle specific challenges related to resolution independence and efficient capturing of multi-scale features in regular grid-based problems, combining the continuous nature of neural operators with the powerful representation capabilities of vision transformers.
@@ 3.1 Architecture Description
The vision transformer encoder takes as input a gridded representation of the input function $u$, yielding a 2+1 dimensional spatio-temporal data tensor $\mathbf{u}\in{\mathbb{R}}^{T\times H\times W\times D}$ with $D$ channels. We patchify our inputs into 3D tokens $\mathbf{u}_{p}\in{\mathbb{R}}^{T\times\frac{H}{P}\times\frac{W}{P}\times C}$ by tokenizing each 2D spatial frame independently, following the process used in standard Vision Transformers (Dosovitskiy et al., 2020). We then add trainable 1D temporal and 2D spatial positional embeddings to each token:
$\mathbf{u}_{pe}=\mathbf{u}_{p}+\text{PE}_{t}+\text{PE}_{s},\quad\text{PE}_{t}\in{\mathbb{R}}^{T\times 1\times 1\times C},\quad\text{PE}_{s}\in{\mathbb{R}}^{1\times\frac{H}{P}\times\frac{W}{P}\times C}.$
@@
To reduce computational cost, we use a temporal aggregation layer based on the Perceiver architecture (Jaegle et al., 2021; Alayrac et al., 2022). Specifically, we learn a predefined number of latent input queries $\mathbf{z}\in{\mathbb{R}}^{1\times C}$ by applying a cross-attention module that maps the flattened visual features $\mathbf{u}_{f}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T\times C}$ to a latent tensor $\mathbf{z}_{agg}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times 1\times C}$ as
$\mathbf{z}^{\prime}$ $=\hat{\mathbf{z}}+\operatorname{MHA}\left(\operatorname{LN}\left(\hat{\mathbf{z}}\right),\operatorname{LN}\left(\mathbf{u}_{f}\right),\operatorname{LN}\left(\mathbf{u}_{f}\right)\right),$
$\mathbf{z}_{agg}$ $=\mathbf{z}^{\prime}+\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}^{\prime}\right)\right).$
Here, $\mathbf{z}$ is initialized by a unit Gaussian distribution and $\hat{\mathbf{z}}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T\times C}$ is obtained by tiling the latent query $\mathbf{z}$ $\left(\frac{H}{P}\times\frac{W}{P}\right)$ times. Besides compressing tokens over the time axis, this module also enables the model to handle inputs with a variable number of time steps.
@@
$\mathbf{z}_{0}$ $=\operatorname{LN}(\mathbf{z}_{agg}),$
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\ell=1\ldots L,$
$\mathbf{z}_{\ell},$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}_{\ell}^{\prime}\right)\right)+\mathbf{z}_{\ell}^{\prime},$ $\ell=1\ldots L.$
@@
$\mathbf{x}^{\prime}_{k}$ $=\mathbf{x}_{k-1}+\operatorname{MHA}\left(\operatorname{LN}\left(\mathbf{x}_{k-1}\right),\operatorname{LN}\left(\mathbf{z}_{L}\right),\operatorname{LN}\left(\mathbf{z}_{L}\right)\right),$ $k=1\ldots K,$
$\mathbf{x}_{k}$ $=\mathbf{x}^{\prime}_{k}+\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{x}^{\prime}_{k}\right)\right),$ $k=1\ldots K.$
@@ 4.1 Main Results
Navier-Stokes equations.
Our third benchmark corresponds to predicting the evolution of a 2D buoyancy-driven flow, depicting a smoke volume rising in a bounded square domain. The underlying governing equations are the incompressible Navier-Stokes equations coupled with a passive scalar transport equation. For a fair comparison with various neural operators and other state-of-the-art deep learning models, we use the dataset generated by PDEArena (Gupta & Brandstetter, 2022), and follow the problem setup reported in Hao et al. (Hao et al., 2024).
Our main results are summarized in Table 3, indicating that CViT outperforms all baselines we have considered. Notably, the previous best result (DPOT-L (Fine-tuned)) is achieved with a much larger transformer-based architecture, which involves pretraining on massive diverse PDE datasets followed by fine-tuning on the target dataset for 500 epochs. Here we also note that DPOT was pretrained and fine-tuned using a rollout loss, which directly optimizes for the multi-step prediction task on which the models are evaluated. In contrast, CViT was trained using a one-step loss without any rollout finetuning. This difference in training objective significantly favors DPOT for rollout evaluations across all benchmarks. Furthermore, CViT exhibits remarkable parameter efficiency. CViT-S, with only 13M parameters, achieves accuracy comparable to the largest pretrained model (1.03B). Additionally, CViT shows good scalability; as the number of parameters increases, performance improves. These findings strongly support the effectiveness our approach in learning evolution operators for complex physical systems. Additional results and sample visualizations are provided in Figure 5, as well as Figure 11 and 12 in Appendix.
Table 3: Relative $L^{2}$ error of rollout predictions on the Navier-Stokes and Diffusion-Reaction benchmark.
Model # Params Navier-Stokes Diffusion-Reaction
FNO 0.5 M 9.12 % 12.00 %
FFNO 1.3 M 8.39 % 5.71 %
GK-T 1.6 M 9.52% 3.59 %
GNOT 1.8 M 17.20 % 3.11 %
Oformer 1.9 M 13.50 % 1.92 %
DPOT-Ti 7 M 12.50 % 3.21 %
MPP-S 30M - 1.12 %
DPOT-S 30 M 9.91 % 3.79 %
MPP-L (Pre-trained) 400 M - 0.98 %
DPOT-L (Pre-trained) 500 M 7.98 % 2.32 %
DPOT-L (Fine-tuned) 500 M 2.78 % 0.73 %
DPOT-H (Pre-trained) 1.03 B 3.79 % 1.91 %
CViT-S 13 M 3.75 % 1.13 %
CViT-B 30 M 3.18 % 1.11 %
CViT-L 92 M 2.35 % 0.68 %
Figure 5: Navier-Stokes benchmark. Representative CViT rollout predictions of the passive scalar field, and point-wise error against the ground truth.
Diffusion-Reaction equations.
@@
Table 3 presents our main results. Consistent with previous findings, CViT-L achieves the best performance across all baselines. CViT models outperform other approaches with comparable or fewer parameters, demonstrating good scalability and parameter efficiency. Representative rollout predictions are provided in Figure 6 and 13 (see Appendix). Our predictions show excellent agreement with the corresponding ground truth.
@@ 5 Discussion
Acknowledgements
We would like to acknowledge support from the US Department of Energy under the Advanced Scientific Computing Research program (grant DE-SC0024563). We also thank the developers of the software that enabled our research, including JAX Bradbury et al. (2018), Matplotlib Hunter (2007), and NumPy Harris et al. (2020).
@@ Acknowledgements
The code used to carry out the ablation experiments is provided as a supplementary material for this submission. If the paper is accepted, we plan on making our entire code-base publicly available on GitHub.
@@ F.2 Model Details
For the shallow water equation and Navier-Stokes equation, the CViT configurations are detailed in Table 1. While we use a patch size of $8\times 8$ by default, we use a smaller patch size of $4\times 4$ when training CViT-L for the Navier-Stokes problem. We find that using a smaller patch size better captures the small-scale features in flow motion, leading to improved accuracy.
@@ F.4 Shallow-Water equations
For training CViT models, please refer to section F.1. For training other baselines in Table LABEL:tab:_swe, we follow the recommended training procedure and hyper-parameters presented in PDEArena (Gupta & Brandstetter, 2022) Specifically, we use the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-5}$. Training is conducted for 50 epochs, minimizing the summed mean squared error. We use cosine annealing as the learning rate scheduler (Loshchilov & Hutter, 2016) with a linear warm-up phase of 5 epochs. An effective batch size of 32 is used for training.
@@
To validate this, we conducted experiments with three different grid resolutions: 95×191,97×95$\times$ 191,97$\times$95 × 191 , 97 × 193 , and $128\times 200$. We performed these experiments using the shallow water equation example under the same hyper-parameter settings. The resulting relative $L^{2}$ errors are summarized below:
@@
Relative $L^{2}$ Error $1.563\times 10^{-2}$ $1.572\times 10^{-2}$ $1.568\times 10^{-2}$ $1.562\times 10^{-2}$
@@
Figure 9: Shallow water benchmark. Ablation studies on the latent dimension of grid features and number of attention heads in decoder.
@@
F.5 Navier-Stokes equation

### v3

@@ Abstract
Operator learning, which aims to approximate maps between infinite-dimensional function spaces, is an important area in scientific machine learning with applications across various physical domains. Here we introduce the Continuous Vision Transformer (CViT), a novel neural operator architecture that leverages advances in computer vision to address challenges in learning complex physical systems. CViT combines a vision transformer encoder, a novel grid-based coordinate embedding, and a query-wise cross-attention mechanism to effectively capture multi-scale dependencies. This design allows for flexible output representations and consistent evaluation at arbitrary resolutions. We demonstrate CViT’s effectiveness across a diverse range of partial differential equation (PDE) systems, including fluid dynamics, climate modeling, and reaction-diffusion processes. Our comprehensive experiments show that CViT achieves state-of-the-art performance on multiple benchmarks, often surpassing larger foundation models, even without extensive pretraining and roll-out fine-tuning. Taken together, CViT exhibits robust handling of discontinuous solutions, multi-scale features, and intricate spatio-temporal dynamics. Our contributions can be viewed as a significant step towards adapting advanced computer vision architectures for building more flexible and accurate machine learning models in the physical sciences. All data and code are publicly available at https://github.com/PredictiveIntelligenceLab/cvit.
@@ 2 Background and Related Work
The emergence of neural operators has driven advancements in solving partial differential equations (PDEs) and modeling physical systems. Building upon the foundational architectures mentioned in the introduction, recent developments have focused on enhancing the expressiveness and efficiency of these models. For instance, extensions of the Fourier Neural Operator (FNO) (Li et al., 2021) have explored factorized representations (Tran et al., 2021) to reduce computational complexity while maintaining performance. Similarly, wavelet-based approaches (Gupta et al., 2021; Tripura & Chakraborty, 2022) have been proposed to capture multi-scale features more effectively. Most existing neural operator architectures can be interpreted as conditioned neural fields (Xie et al., 2022) with specific choices of base fields and conditioning mechanisms, see Appendices B and C for a detailed discussion.
Transformer-Based Operator Learning.
@@
Concurrently, the success of vision transformers in computer vision tasks has inspired their adaptation to scientific machine learning problems. The self-attention mechanism central to transformers offers a powerful tool for capturing long-range dependencies in spatial and temporal data, a crucial aspect in many PDE-governed systems. For example, OFormer (Li et al., 2022) introduced a novel way to embed continuous input functions and output queries into a transformer architecture. This approach demonstrated the potential of transformers in handling the inherent continuity of physical problems. GNOT (Hao et al., 2023) further extended this idea by incorporating graph structures, allowing for more flexible representation of input functions and improved handling of irregular domains. A significant advancement in this direction came with DPOT (Hao et al., 2024), which introduced a denoising pre-training strategy. This approach enables the model to learn from diverse PDE datasets, significantly enhancing its generalization capabilities.
Open Challenges.
@@
Despite these advancements, several key challenges remain in the field of operator learning, particularly in addressing resolution independence, high-dimensional data, long-range spatio-temporal dependencies and trade-offs between model expressiveness and computational efficiency. Transformers typically rely on fixed-resolution data and positional encodings, which hinder their ability to generalize across varying scales or irregular domains commonly encountered in scientific problems. Efficiently capturing global and local dependencies in high-dimensional systems further compounds computational demands, especially as self-attention scales quadratically with input size. These constraints make it difficult to balance scalability, precision, and efficiency. Addressing even a subset of these challenges requires novel architectural designs that can leverage the strengths of both neural operators and vision transformers while overcoming their respective limitations. This sets the stage for our proposed Continuous Vision Transformer (CViT), which aims to tackle specific challenges related to resolution independence and efficient capturing of multi-scale features in regular grid-based problems, combining the continuous nature of neural operators with the powerful representation capabilities of vision transformers.
@@ 3.1 Architecture Description
The vision transformer encoder takes as input a gridded representation of the input function $u$, yielding a spatio-temporal data tensor $\mathbf{u}\in{\mathbb{R}}^{T\times H\times W\times D}$ with $D$ channels. We patchify our inputs into 3D tokens $\mathbf{u}_{p}\in{\mathbb{R}}^{T\times\frac{H}{P}\times\frac{W}{P}\times C}$ by tokenizing each 2D spatial frame independently, following the process used in standard Vision Transformers (Dosovitskiy et al., 2020). We then add trainable 1D temporal and 2D spatial positional embeddings to each token:
$\mathbf{u}_{pe}=\mathbf{u}_{p}+\text{PE}_{t}+\text{PE}_{s},\quad\text{PE}_{t}\in{\mathbb{R}}^{T\times 1\times 1\times C},\quad\text{PE}_{s}\in{\mathbb{R}}^{1\times\frac{H}{P}\times\frac{W}{P}\times C}.$ (1)
@@
To reduce computational cost, we use a temporal aggregation layer based on the Perceiver architecture (Jaegle et al., 2021; Alayrac et al., 2022). Specifically, we learn a pre-defined number of latent input queries $\mathbf{z}\in{\mathbb{R}}^{T^{\prime}\times C}$. These queries serve as learnable parameters in a cross-attention module, which processes the visual features of our input. The Perceiver module operates on flattened visual features $\mathbf{u}_{f}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T\times C}$ obtained by flattening the positionally encoded features $u_{pe}$. Through this cross-attention mechanism, we aggregate the temporal information into a compact latent representation $\mathbf{z}_{agg}\in\mathbb{R}^{(\frac{H}{P}\times\frac{W}{P})\times T^{\prime}\times C}$ as:
$\mathbf{z}^{\prime}$ $=\hat{\mathbf{z}}+\operatorname{MHA}\left(\operatorname{LN}\left(\hat{\mathbf{z}}\right),\operatorname{LN}\left(\mathbf{u}_{f}\right),\operatorname{LN}\left(\mathbf{u}_{f}\right)\right),$ (2)
$\mathbf{z}_{agg}$ $=\mathbf{z}^{\prime}+\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}^{\prime}\right)\right).$ (3)
Here, $\mathbf{z}$ is initialized by a unit Gaussian distribution and $\hat{\mathbf{z}}\in{\mathbb{R}}^{(\frac{H}{P}\times\frac{W}{P})\times T^{\prime}\times C}$ is obtained by tiling the latent query $\mathbf{z}$ $\left(\frac{H}{P}\times\frac{W}{P}\right)$ times. Besides compressing tokens over the time axis, this module also enables the model to handle inputs with a variable number of time steps.
@@
$\mathbf{z}_{0}$ $=\operatorname{LN}(\mathbf{z}_{agg}),$ (4)
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\ell=1\ldots L,$ (5)
$\mathbf{z}_{\ell},$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}_{\ell}^{\prime}\right)\right)+\mathbf{z}_{\ell}^{\prime},$ $\ell=1\ldots L.$ (6)
@@
$\mathbf{x}^{\prime}_{k}$ $=\mathbf{x}_{k-1}+\operatorname{MHA}\left(\operatorname{LN}\left(\mathbf{x}_{k-1}\right),\operatorname{LN}\left(\mathbf{z}_{L}\right),\operatorname{LN}\left(\mathbf{z}_{L}\right)\right),$ $k=1\ldots K,$ (8)
$\mathbf{x}_{k}$ $=\mathbf{x}^{\prime}_{k}+\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{x}^{\prime}_{k}\right)\right),$ $k=1\ldots K.$ (9)
@@ 4.1 Main Results
Navier-Stokes equations (NS).
Our third benchmark focuses on the compressible and incompressible Navier-Stokes equations, which underpin a variety of systems, ranging from hydromechanical processes to weather forecasting and turbulent dynamics analysis. For a fair comparison with various neural operators and other state-of-the-art deep learning models, we use the dataset generated by PDEArena (Gupta & Brandstetter, 2022), and PDEBench Takamoto et al. (2022). and follow the problem setup reported in Hao et al. (Hao et al., 2024).
Our main results are summarized in Table 3, indicating that CViT outperforms all baselines we have considered. Notably, the previous best result (DPOT-L (Fine-tuned)) is achieved with a much larger transformer-based architecture, which involves pretraining on massive diverse PDE datasets followed by fine-tuning on the target dataset for 500 epochs. Here we also note that DPOT was pretrained and fine-tuned using a rollout loss, which directly optimizes for the multi-step prediction task on which the models are evaluated. In contrast, CViT was trained using a one-step loss without any rollout finetuning. This difference in training objective significantly favors DPOT for rollout evaluations across all benchmarks. Furthermore, CViT exhibits remarkable parameter efficiency. CViT-S, with only 13M parameters, achieves accuracy comparable to the largest pretrained model (1.03B). Additionally, CViT shows good scalability; as the number of parameters increases, performance improves. These findings strongly support the effectiveness our approach in learning evolution operators for complex physical systems. Additional results and sample visualizations are provided in Figure 5, as well as Figure 13 and 14 in Appendix.
Table 3: Relative $L^{2}$ error of rollout predictions on the incompressible and compressible Navier-Stokes and Diffusion-Reaction benchmark.
Model # Params NS CNS DR
FNO 0.5 M 9.12 % 9.60 % 12.00 %
FFNO 1.3 M 8.39 % 5.20 % 5.71 %
GK-T 1.6 M 9.52% 3.77% 3.59 %
GNOT 1.8 M 17.20 % 4.20 % 3.11 %
Oformer 1.9 M 13.50 % 6.25 % 1.92 %
DPOT-Ti 7 M 12.50 % 3.97 % 3.21 %
MPP-S 30M - - 1.12 %
DPOT-S 30 M 9.91 % 3.37 % 3.79 %
MPP-L (Pre-trained) 400 M - - 0.98 %
DPOT-L (Pre-trained) 500 M 7.98 % 2.16 % 2.32 %
DPOT-L (Fine-tuned) 500 M 2.78 % 1.31 % 0.73 %
DPOT-H (Pre-trained) 1.03 B 3.79 % 1.80 % 1.91 %
CViT-S 13 M 3.75 % 2.71 % 1.13 %
CViT-B 30 M 3.18 % 1.99 % 1.11 %
CViT-L 92 M 2.35 % 1.29 % 0.68 %
Figure 5: Incompressible Navier-Stokes benchmark (NS). Representative CViT rollout predictions of the passive scalar field, and point-wise error against the ground truth.
Figure 6: Compressible Navier-Stokes Benchmark (CNS). Representative CViT rollout prediction of the density field $\rho$, and point-wise error against the ground truth.
Diffusion-Reaction equations (DR).
@@
Table 3 presents our main results. Consistent with previous findings, CViT-L achieves the best performance across all baselines. CViT models outperform other approaches with comparable or fewer parameters, demonstrating good scalability and parameter efficiency. Representative rollout predictions are provided in Figure 18 and 19 (see Appendix). Our predictions show excellent agreement with the corresponding ground truth.
@@ 5 Discussion
Acknowledgment
We would like to acknowledge support from the US Department of Energy under the Advanced Scientific Computing Research program (grant DE-SC0024563) and Yale Institute for Foundations of Data Science. We also thank the developers of the software that enabled our research, including JAX Bradbury et al. (2018), Matplotlib Hunter (2007), and NumPy Harris et al. (2020).
@@ Acknowledgment
The code used to carry out the ablation experiments is provided as a supplementary material for this submission. All data and code are publicly available at https://github.com/PredictiveIntelligenceLab/cvit.
@@ F.2 Model Details
For the shallow water equation, compressible and incompressible Navier-Stokes equations and diffusion reaction equations, the CViT configurations are detailed in Table 1. While we use a patch size of $8\times 8$ by default, we use a smaller patch size of $4\times 4$ when training CViT-L for the Navier-Stokes problem. We find that using a smaller patch size better captures the small-scale features in flow motion, leading to improved accuracy.
@@ F.4 Shallow-Water equations
For training CViT models, please refer to section F.1. For training other baselines in Table 2, we follow the recommended training procedure and hyper-parameters presented in PDEArena (Gupta & Brandstetter, 2022) Specifically, we use the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-5}$. Training is conducted for 50 epochs, minimizing the summed mean squared error. We use cosine annealing as the learning rate scheduler (Loshchilov & Hutter, 2016) with a linear warm-up phase of 5 epochs. An effective batch size of 32 is used for training.
@@
To validate this, we use CViT-B as our backbone architecture, conduct experiments with three different grid resolutions: 95×191,97×95$\times$ 191,97$\times$95 × 191 , 97 × 193 , and $128\times 200$. We performed these experiments using the shallow water equation example under the same hyper-parameter settings. The resulting relative $L^{2}$ errors are summarized below:
@@
Rel. $L^{2}$ Error $1.563\times 10^{-2}$ $1.572\times 10^{-2}$ $1.568\times 10^{-2}$ $1.562\times 10^{-2}$
@@
Figure 9: Shallow water benchmark. Ablation studies on the latent dimension of grid features and number of attention heads in decoder. Here we use CViT-B as our backbone architecture.
Impact of random Fourier feature frequencies.
We conduct an ablation study to investigate the impact of the scaling parameter $s$ in random Fourier feature embeddings. Using CViT-B as our backbone architecture, we trained models with different values of $s$ on the SWE benchmark. As shown in Table 7, model accuracy improves with moderately large values of $s$ but deteriorates when $s$ becomes too large (e.g., $s=16\pi$), reverting to performance levels similar to $s=4\pi$. Notably, our proposed grid-based embedding consistently outperforms the random Fourier feature approach across all tested values of $s$.
Table 7: Shallow water benchmark. Ablation studies on the hyper-parameter $s$ of random Fourier features. Here we use CViT-B as the backone.
Random Fourier Features Grid-based embedding
$s=2\pi$ $s=4\pi$ $s=8\pi$ $s=16\pi$ $\beta=10^{5}$
Rel. $L^{2}$ error $4.94\times 10^{-2}$ $4.16\times 10^{-2}$ $3.25\times 10^{-2}$ $4.08\times 10^{-2}$ $2.69\times 10^{-2}$
@@
F.5 Incompressible Navier-Stokes equation